The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: Eva K. Grebel, You-Hua Chu, and Dominik J. Bomans

                                 mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 16 December 29, 1997

Best wishes for a happy and successful year 1998!

Table of Contents

Abstracts of 13 refereed papers 1
Thesis abstract  9
Meeting announcements  10

Abstracts of Refereed Papers

Molecular Cloud Structure in the Magellanic Clouds: Effect of Metallicity

Soojong Pak^{1,2}, D.T. Jaffe^{1}, Ewine F. van Dishoeck^{3}, L.E.B. Johansson^{4}, & R.S. Booth^{4}

1 The University of Texas at Austin, Department of Astronomy, Austin, TX 78712, USA 2 Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, 85748 Garching, Germany 3 Sterrewacht Leiden, P.O.Box 9513, 2300 RA Leiden, The Netherlands 4 Onsala Space Observatory, S-439 92 Onsala, Sweden

The chemical structure of neutral clouds in low metallicity environments is examined with particular emphasis on the H to H_{2} and C^{+} to CO transitions. We observed near-IR H_{2} lines and the CO J=1-0 line from 30 Doradus and N159/N160 in the Large Magellanic Cloud and from DEM S 16, DEM S 37, and LI-SMC 36 in the Small Magellanic Cloud. We find that the H_{2} emission is UV-excited and that (weak) CO emission always exists (in our surveyed regions) toward positions where H_{2} and [CII] emission have been detected. Using a PDR code and a radiative transfer code, we simulate the emission of line radiation from spherical clouds and from large planar clouds. Because the [CII] emission and H_{2} emission arise on the surface of the cloud and the lines are optically thin, these lines are not affected by changes in the relative sizes of the neutral cloud and the CO bearing core, while the optically thick CO emission can be strongly affected. The sizes of clouds are estimated by measuring the deviation of CO emission strength from that predicted by a planar cloud model of a given size. The average cloud column density and therefore size increases as the metallicity decreases. Our result agrees with the photoionization regulated star formation theory by McKee (1989).

Accepted by: The Astrophysical Journal

For preprints, contact soojong@mpe-garching.mpg.de
Also available from the URL http://mpe3d.mpe-garching.mpg.de/~soojong
   and http://xxx.lanl.gov/abs/astro-ph/9712158
or by anonymous ftp at mpe3d.mpe-garching.mpg.de/pub/soojong/mcs/pak98.ps.gz

On the Form of the H II Region Luminosity Function

M.S. Oey and C.J. Clarke

Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, U.K.

Observed variations in the H II region luminosity function (H II LF) seen in spiral arm vs. interarm regions, and different galactic Hubble type, can be explained simply by evolutionary effects and maximum number of ionizing stars per cluster. We present Monte Carlo simulations of the H II LF, drawing the number of ionizing stars N_{∗} from a power-law distribution of constant slope, and the stellar masses from a Salpeter IMF with an upper-mass limit of 100 M_{⊙}. We investigate the evolution of the H II LF, as determined by stellar main-sequence lifetimes and ionizing luminosities, for a single burst case and continuous creation of the nebular population. Shallower H II LF slopes measured for the arms of spiral galaxies can be explained as a composite slope, expected for a zero-age burst population, whereas the interarm regions tend to be dominated by evolved rich clusters described by a single, steeper slope. Steeper slopes in earlier-type galaxies can be explained simply by a lower maximum N_{∗} cutoff found for the parent OB associations. The form of the H II LF can reveal features of the most recent (∼ < 10 Myr) star formation history in nearby galaxies.

Accepted by: The Astronomical Journal For preprints, contact oey@ast.cam.ac.uk

Also available from the URL http://www.ast.cam.ac.uk/~oey/oeypubs.html
   or http://xxx.lanl.gov/abs/astro-ph/9712157

A radio continuum study of the Magellanic Clouds. Part VII: Discrete radio sources in the Magellanic Clouds

M.D. Filipović^{1,2,3}, R.F. Haynes^{3,1}, G.L. White^{1}, and P.A. Jones^{1}

1 University of Western Sydney, Nepean, P.O. Box 10, Kingswood, NSW 2747 Australia 2 Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, D-85740 Garching, Germany 3 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121 Australia

We present a study of discrete radio sources in the Magellanic Clouds (MCs) using the latest largescale radio surveys made with the Parkes radio telescope between 1.4 and 8.55 GHz. These surveys achieved higher sensitivity than previous surveys done with the Parkes telescope and so the number of discrete radio sources detected towards the MCs has increased by factor of five. Also, we have obtained improved positions, flux densities, and radio spectral indices for all of these sources.

A total of 483 sources towards the Large Magellanic Cloud (LMC) and 224 towards the Small Magellanic Cloud (SMC) have been detected at at least one radio frequency. Most of the MC’s sources have been classified in one of three groups: SNRs, H ii regions, or background sources according to classification criteria established here. In total, 209 discrete radio sources in the LMC and the 37 sources in the SMC are classified here to be either H ii regions or SNRs. We investigate their luminosity functions as well as the statistics of background sources behind the MCs. Also, we examine the distribution of SNRs and H ii regions in the MCs.

Accepted by: Astronomy & Astrophysics Supplement

For preprints, contact m.filipovic@uws.edu.au

A radio continuum study of the Magellanic Clouds. Part VIII: Discrete sources common to radio and infrared surveys

of the Magellanic Clouds

M.D. Filipović^{1,2,3}, P.A. Jones^{1}, G.L. White^{1}, and R.F. Haynes^{3,1}

1 University of Western Sydney, Nepean, P.O. Box 10, Kingswood, NSW 2747 Australia 2 Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, D-85740 Garching, Germany 3 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121 Australia

We compare Parkes Telescope radio surveys with the IRAS Infrared (IR) surveys of the Magellanic Clouds (MCs). We find 130 discrete sources in common towards the Large Magellanic Cloud (LMC) with both radio and IR emission. These 130 sources are mainly H ii regions (89) and supernova remnants (21). For 12 of the sources we have no identification and eight are background objects. We find 38 sources in common for the Small Magellanic Cloud (SMC). Most of these sources are intrinsic (31) to the SMC, five sources are previously known background galaxies, and two sources remain ambiguous.

A flux density comparison of the radio and IR sources shows very good correlation and we note that the strongest sources at both radio and IR frequencies are H ii regions. From the radio–IR comparison we propose that some 40 new sources in the LMC and 10 in the SMC are H ii regions or SNRs. All these new sources are also identified in optical surveys.

Submitted to: Astronomy & Astrophysics Supplement

For preprints, contact m.filipovic@uws.edu.au

Five Mature Supernova Remnants in the Large Magellanic Cloud

John R. Dickel^{1} and D.K. Milne^{2}

1 Astronomy Department, University of Illinois, 1002 West Green Street, Urbana IL 61801, USA 2 Australia Telescope National Facility, CSIRO, Box 76, Epping NSW 2121, Australia

Radio observations of the extended supernova remnants N23:SNR, N120:SNR, 0525-660, N49, and N44:Shell 3 in the Large Magellanic Cloud show a variety of structures and little difference from shell SNRs in the Milky Way. N23:SNR and N120:SNR have very regular outlines but the shells are much brighter on the sides toward the most prominent parts of their associated Hii region complexes, probably reflecting density gradients in the surroundings. N23:SNR has a mixed magnetic field orientation and a mean magnetic field strength of 15 microGauss. Its magnetic energy is about 4 times less than the relativistic electron energy. 0525-660 appears to be isolated in the northern region of the LMC; it has a partial shell of thin filaments and some bright, patchy emission across the face of the remnant. N49, one of the brightest SNRs in the LMC, shows the strongest polarized emission in the west where the total intensity is very weak, probably as a result of high magnetic field organization. N44:Shell 3 on the northeastern edge of a very large Hii region complex is definitely confirmed as a SNR but has a different central position and outline at radio wavelengths than it does at X-ray wavelengths.

Accepted by: The Astronomical Journal

For preprints, contact johnd@astro.uiuc.edu

Mass-Loss Rate and He/H Abundance of the Erupting Component in the SMC system HD 5980

G. Koenigsberger ^{1,3}, M. Peña^{1,4}, W. Schmutz^{2}, and S. Ayala^{1}

1 Instituto de Astronomia, UNAM, Apdo. Postal 70-264, Mexico D.F. 04510 2 Institute of Astronomy, ETH Zürich, CH-8092 Zürich, Switzerland 3 Guest Observer with the IUE satellite, which is sponsored and operated by NASA, ESA and the Science and Engineering Council of the UK 4 Visiting Astronomer at Cerro Tololo Inter-American Observatory, which is operated by the Association of Universities for Research in Astronomy, Inc. under cooperative agreement with the NSF

The binary Wolf-Rayet system HD 5980 in the Small Magellanic Cloud underwent a major LBVtype eruptive event in 1994. It is the first such recorded event in which the apparent precursor transits from WNE to WNL spectral types prior to the eruption. In this paper we analyze the spectrum of the system obtained at the time when the outburst was declining (1994 December), but when the dominant spectrum was that of the eruptor. From the non-LTE analysis we obtain Ṁ = 10^{−3}M_{⊙} yr^{−1}, N[He]/N[H] = 0.43 (by number; Y =0.63 by mass), v_{∞} = 600 km s^{−1}, T_{∗} = 35 500 K, and L_{eruptor} = 3 × 10^{6} L_{⊙}. A comparison of the He/H abundance derived here and an estimate obtained from published data of 20 years ago leads us to conclude that the member of the binary system which underwent the eruption is the star formerly classified as an O7 supergiant. The considerable amount of He in this star indicates that it is in the transition to becoming an H-poor WR. By comparing the stellar parameters with single star evolutionary tracks we derive that the progenitor was more massive than 120 M_{⊙}, and its current mass is close to 80 M_{⊙}.

Accepted by: The Astrophysical Journal For preprints, contact G. Koenigsberger, M. Peña: gloria@astroscu.unam.mxh

Also available by anonymous ftp at kepler.astroscu.unam.mx, cd incoming/koenigsberger/paper2

Wind Velocity Variations in the LBV-Type Erupting Star of the Wolf-Rayet Binary HD 5980

G. Koenigsberger ^{1,4}, L.H. Auer^{2}, L. Georgiev^{1}, and E. Guinan^{3,4}

1 Instituto de Astronomia, UNAM, Apdo. Postal 70-264, Mexico D.F. 04510 2 Los Alamos National Laboratory, P.O. Box 1663, Los Alamos, NM, USA 3 Department of Astronomy and Astrophysics, Villanova University, Villanova, PA 19085, USA 4 Guest Observer with the IUE satellite, which is sponsored and operated by NASA, ESA and the Science and Engineering Council of the UK

We present the wind velocity and UV luminosity variations in the Wolf-Rayet system HD 5980 obtained over a time span during which one of the stars of the system was transformed into a Luminous Blue Variable and underwent an eruption. We are able to separate the velocity components of the two stars in the system: A stable velocity component at −1700 km s^{−1} is associated with the non-erupting star, while the variable wind with velocities ranging from −500 to − 3000 km s^{−1} corresponds to the eruptor. The development of a fast wind following the slow wind eruptive phase is observed. Under the assumption of radiatively driven winds, these changing velocities indicate that the radius of the photosphere gradually increased during at least 12 years prior to the 1994 eruption, decreasing rapidly thereafter. An estimate of the stellar parameters indicates that the erupting star is massive (M > 40 M_{⊙}) and very luminous (L > 10^{6} L_{⊙}), and that during the eruption its radius extended beyond the binary orbit (R_{∗} > 100 R_{⊙}).

Accepted by: The Astrophysical Journal For preprints, contact G. Koenigsberger gloria@astroscu.unam.mx

Also available by anonymous ftp at kepler.astroscu.unam.mx , cd /incoming/koenigsberger/paper1

The Lutz-Kelker Bias in Trigonometric Parallaxes

René D. Oudmaijer^{1}, Martin A.T. Groenewegen^{2}, and Hans Schrijver^{3}

1Blackett Laboratory, Imperial College of Science,Technology & Medicine,Prince Consort Road, London SW7 2BZ, U.K. 2Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Straße 1, D-85740 Garching, Germany 3SRON, Sorbonnelaan 2, NL-3584 CA Utrecht, The Netherlands

The theoretical prediction that trigonometric parallaxes suffer from a statistical effect has become topical again now that the results of the Hipparcos satellite have become available. This statistical effect, the so-called Lutz-Kelker bias, causes measured parallaxes to be too large. This has the implication that inferred distances and hence inferred luminosities are too small. Published analytic calculations of the Lutz-Kelker bias indicate that the inferred luminosity of an object is, on average, 30% too small when the error in the parallax is only 17.5%. Yet, this bias has never been determined empirically. In this paper we investigate whether there is such a bias by comparing the best Hipparcos parallaxes with ground-based measurements. We find that there is indeed a large bias affecting parallaxes, with an average and scatter comparable to predictions. We propose a simple method to correct for the LK bias, and apply it successfully to a sub-sample of our stars. We then analyze the sample of 26 ‘best’ Cepheids used by Feast & Catchpole (1997) to derive the zero-point of the period-luminosity relation. The final result is based on the 20 fundamental mode pulsators and leads to a distance modulus to the Large Magellanic Cloud - based on Cepheid parallaxes- of 18.56 ± 0.08, consistent with previous estimates.

Accepted by: MNRAS Letters

For preprints, contact r.oudmaijer@ic.ac.uk

The Young Inter-Cloud Populations. II. The Midwest of the Large Magellanic Cloud

P. Battinelli^{1} and S. Demers^{2}

1 Osservatorio Astronomico di Roma, Viale del Parco Mellini 84 I-00136 Rome, Italy 2 Département de Physique, Université de Montréal, Montreal, Canada H3C 3J7

We present the BV CCD photometry of a field of blue stars in the midwest of the Large Magellanic Cloud (LMC), located nearly three degrees from its center. This stellar population, first noticed by Pierre [A&A, 175, 54, 1987], is found to consist of stars of all ages larger than ∼ 30 Myr along with a substantial 4.5 Gyr old population. This field corresponds to the outermost known region of blue stars in the southwest quadrant of the LMC. The slope of the observed differential luminosity function (LF) for the main sequence stars is in good agreement with other determinations of LMC field regions but differs significantly from the LF of the inter-Cloud aggregates. We conclude that because this field does not show, in recent time, any significant burst of star formation that it is not chronologically linked to the LMC/SMC close encounter. Therefore, star formation must have taken place continuously in that area for at least ∼ 1 Gyr.

Accepted by: The Astronomical Journal

For preprints, contact battinel@oarhp1.rm.astro.it

Observational discovery of the AGB–bump in densely populated color-magnitude diagrams of galaxies and star clusters

C. Gallart^{1}

1 Carnegie Observatories, 813 Santa Barbara St., Pasadena, CA 91101, USA

The huge photometric databases that are being created for stars in the Magellanic Clouds and in Local Group galaxies not only allow detailed study of the star formation history of these systems, but also provide important information on stellar evolutionary theory. Two low-level features above the red clump in the LMC color-magnitude diagram (CMD) have been discussed by Zaritsky & Lin (1997) and Alcock et al. (1997) as the possible signature of an intervening population. I conclude that one of the features, which has also been associated with the red giant branch bump (RGB–bump) predicted by theory and observed in globular clusters, is instead produced during asymptotic giant branch (AGB) evolution. I will call it the AGB–bump. In this paper, it will be shown that the stellar evolution predictions about the position and strength of the AGB–bump are in very good agreement with the observed structures in the LMC, M31, and globular clusters. The position in the CMD of the other feature is consistent with the location of the blue loops of a few Myr old stars and it has also been observed in the CMD of other galaxies with a young stellar population.

Accepted by: The Astrophysical Journal Letters For preprints, contact carme@ociw.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9712283
or by anonymous ftp at ociw.edu; pub/carme

Two serendipitous low-mass LMC clusters discovered with HST

Basílio X. Santiago^{1}, Rebecca A.W. Elson^{2}, Steinn Sigurdsson^{2}, and Gerard F. Gilmore^{2}

1 Departamento de Astronomia, Universidade Federal do Rio Grande do Sul, 91501-970, Porto Alegre, RS, Brazil 2 Institute of Astronomy, Cambridge University, Madingley Road, Cambridge CB3 0HA, United Kingdom

We present V and I photometry of two open clusters in the LMC down to V ∼ 26. The clusters were imaged with the Wide Field and Planetary Camera - 2 on board of the Hubble Space Telescope, as part of the Medium Deep Survey Key-Project. Both are low luminosity (M_{V} ∼ −3.5), low mass systems (M ∼ 10^{3}M_{⊙}). The chance discovery of these two clusters in two parallel WFPC2 fields suggests a significant incompleteness in the LMC cluster census near the bar. One of the clusters is roughly elliptical and compact, with a steep light profile, a central surface brightness μ_{V} (0) ∼ 20.2 mag/arcsec^{2}, halflight radius r_{hl} ∼ 0.9 pc (total visual major diameter D ∼ 3 pc) and an estimated mass M ∼ 1500M_{⊙}. From the colour-magnitude diagram and isochrone fits we estimate its age as τ ∼ 2 − 5 × 10^{8} years. Its mass function has a fitted slope of Γ = Δlogϕ(M)/ΔlogM = −1.8 ± 0.7 in the range probed (0.9 < ∼ M/M_{⊙} < ∼ 4.5). The other cluster is more irregular and sparser, having shallower density and surface brightness profiles. We obtain Γ = −1.2 ± 0.4, and estimate its mass as M ∼ 400M_{⊙}. A derived upper limit for its age is τ < 5 × 10^{8} years. Both clusters have mass functions with slopes similar to that of R136, a massive LMC cluster, for which HST results indicate Γ ∼ −1.2. They also seem to be relaxed in their cores and well contained in their tidal radii.

Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact santiago@if.ufrgs.br
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9711356

WFPC2 Observations of the Small Magellanic Cloud Intermediate-Age Populous Cluster NGC 416

Kenneth J. Mighell^{1}, Ata Sarajedini^{2}, and Rica S. French^{3}

1Kitt Peak National Observatory, NOAO, P.O. Box 26732, Tucson, AZ 85726-6732, USA 2Department of Physics & Astronomy, San Francisco State Univ.,1600 Holloway Avenue, San Francisco, CA 94132, USA 3Physics & Astronomy Department, WPS 219, P. O. Box 71, Murfreesboro, TN 37132, USA

We present our analysis of archival Hubble Space Telescope Wide Field Planetary Camera 2 observations in F555W (∼V) and F450W (∼B) of the intermediate-age populous star cluster NGC 416 in the Small Magellanic Cloud galaxy. We use published photometry of two other SMC populous star clusters, Lindsay 1 and Lindsay 113, to investigate the age sequence of these three star clusters. We estimate that these clusters have age ratios of age_{NGC416}/age_{L1} ≈ 0.73±0.05 and age_{L113}/age_{L1} ≈ 0.52±0.09 using an extrapolation of the d_{(B−V)} method (which uses the color difference between the red horizontal branch and the red giant branch as an age indicator) of Sarajedini, Lee, & Lee [ApJ, 450, 712 (1995)] . These age ratios provide absolute age estimates of 6.6±0.5 Gyr and 4.7±0.8 Gyr for NGC 416 and Lindsay 113, respectively, assuming that Lindsay 1 is 9 Gyr old. Metallicities of [Fe/H]=−1.44±0.12, −1.35±0.08, −1.24±0.11 dex, and reddenings of E(B−V)=0.08±0.03, 0.06±0.02, 0.00±0.02 mag for NGC 416, Lindsay 1, and Lindsay 113, respectively, were determined using the simultaneous reddening and metallicity (SRM) method of Sarajedini & Layden [AJ, 113, 264, (1997)]. Accurate (relative) ages for the intermediate-age populous clusters in the Small Magellanic Cloud (e.g. via deep main sequence photometry) would allow the d_{(B−V)} method to be recalibrated with star clusters that are significantly younger than 7 Gyr. An extended d_{(B−V)} method could prove to be a very useful age diagnostic for future studies of the intermediate-age metal-poor stellar populations in Local Group galaxies where accurate main-sequence turnoff photometry at M_{V} ≈+4 mag is currently not possible or practical.

Accepted by: Astrophysical Journal Letters For preprints, contact mighell@noao.edu

Also available from the URL http://www.noao.edu/noao/staff/mighell/ngc416
   or http://xxx.lanl.gov/abs/astro-ph/9712314

A Hybrid N-body/Cellular Automaton Scheme for Modelling Propagating Star Formation in Galaxies

L.T. Gardiner, C. Turfus, and M. Wang

Sun Moon University, Asan-kun, Republic of Korea

We present a numerical scheme for modelling star formation in galaxies based on the dual operation of gravitational instabilities and propagating star formation. Star formation is superposed on a mechanical N-Body simulation of a collisional particle system by using a cellular automaton framework. The simulations involve 5000 particles representing gas clouds which can collide inelastically, and have been performed on a 2D square surface with quasiperiodic boundary conditions. The Large Magellanic Cloud (LMC) has been used as a benchmark galaxy for choosing various parameters for the model setup as well as for comparing simulation results. The input of kinetic energy from starforming regions leads to self-regulation of the global dynamics of star formation. The model naturally generates cloud complexes assembled from groups of cloud particles and simulates the ignition and propagation of star formation in such complexes leading to the formation of low density bubbles and expanding shells. The percentage of the model surface undergoing star formation and the ratio of the rates of stimulated to spontaneous star formation are shown to be consistent with models based on the theory of Stochastic Self-Propagating Star Formation (SSPSF). The simulations produce global fluctuations in star-formation rates which suggest the possible existence of large-scale coordinated star-forming activity in disc galaxies. The code is shown to lead to star formation that propagates very nearly isotropically and can thus be adapted to the case of rotating disc galaxies without geometrical modification. It is currently being applied to a direct simulation of the global star formation characteristics of the LMC.

Submitted to: Monthly Notices of the Royal Astronomical Society

For preprints, contact ltg@omega.sunmoon.ac.kr
Also available from the URL http://alpha.sunmoon.ac.kr/astro/

Thesis Abstract

Analysis of Three Fast Oxygen-Neon-Magnesium Novae

Karen M. Vanlandingham

Arizona State University, Dept. of Physics and Astronomy, Tempe, AZ 85287-1504, USA

Since astronomers have only known about the existence of oxygen-magnesium-neon (ONeMg) novae for the last decade or so, little work has been done on them as a group. It is the purpose of this work to study three fast ONeMg novae: V838 Nova Herculis 1991, V693 Nova Corona Austrinae 1981, and Nova LMC 1990#1. Each of these novae have been studied in detail and a comparison of the results is provided in order to better understand ONeMg novae as a class. These analyses have provided values of the elemental abundances, ejection velocities, ejected masses, white dwarf masses, maximum outburst luminosities, and white dwarf turn-off time for each of these novae. The results of this study can be used in many ways to help further the understanding of novae and stellar evolution. The elemental abundances that have been determined can be used as input to theoretical hydrodynamic models. This will provide better constraints on the models in order to determine the energetics of the outburst and the mass of the underlying white dwarf, properties which are difficult to obtain observationally. The determination of the abundances will also help determine the contribution of fast ONeMg novae to the heavy element content of the interstellar medium (ISM). In particular, with the abundance results, the mass fraction of ^{26}Al and ^{22}Na in the novae ejecta can be estimated. These estimates will help to determine if novae are important for galactic nucleosynthesis. Analysis of Nova LMC 1990#1 provides the means to compare Galactic with extragalactic novae. The results for Nova LMC 1990#1, combined with results from other Large Magellanic Cloud novae, can be used to calibrate the luminosities of (and therefore distances to) Galactic novae. This calibration may allow a better determination of the fast end of the absolute-magnitude decay-rate relationship (or perhaps develop a separate relationship for these objects). A tight, well-defined relationship is crucial for novae to be reliable standard candles.

Ph.D. obtained from Arizona State University on December 11, 1997 under direction of Prof. Sumner Starrfield.

For copies or related preprints, contact karenv@suelee.la.asu.edu

Also available from the URL http://brian.la.asu.edu/~karenv/research.html
or by anonymous ftp at ftp://calvin.physast.uga.edu/pub/preprints/

Meeting Announcements

3rd Cologne-Zermatt Symposium The Physics and Chemistry of the Interstellar Medium

Date and venue Following the meetings in 1988 and 1993 the 3rd Cologne-Zermatt Symposium “The Physics and Chemistry of the Interstellar Medium” will be held in Zermatt, September 22nd - 25th, 1998.

Scope of the colloquium We hope to address a substantial fraction of the astrophysics community active in interstellar medium research. With the recent exciting results from ISO and new advances in observing techniques, instrumentation, and theoretical modelling we foresee exciting presentations and a stimulating exchange of ideas.

Over four days, the symposium will include invited review talks along with contributed talks and poster presentations.

Outline of the scientific program The major topics are:

• Star Formation in the Early Universe

• The ISM in External Galaxies

• Structure of the Milky Way

• Structure and Evolution of the ISM

• The ISM in Star-Forming Regions

• Chemistry, Photochemistry, and Laboratory Astrophysics

• Observatories, Telescopes, and Instrumentation

Scientific Organizing Committee The scientific organizing committee responsible for the scientific scope of the symposium consists of: L. Blitz, R. Booth, E. van Dishoeck, R. Genzel, M. Guélin, R. Kawabe, J. Martin-Pintado, K. Menten, P. Myers, T. Phillips, D.A. Williams.

Web page For further information and registration please visit the conference webpage at

http://www.ph1.uni-koeln.de/zermatt1998/ . Registration deadline is April 15, 1998.

Local Organizing Committee In case of additional questions, you can contact the local organizing committee at Cologne (G. Winnewisser, J. Stutzki, C. Kramer, V. Ossenkopf, U. Corneliussen). You can reach us by email at zermatt@ph1.uni-koeln.de, via telephone at +49 221 470 3554, or by fax at +49 221 470 5162.

Colloquium of Paris-Meudon Observatory Galaxy evolution: Connecting the distant Universe with the Local Fossil Record

Date and venue The colloquium will be held at the Paris Observatory at Meudon, France, on September 21-25, 1998. It will be held in honour of Roger and Giusa Cayrel, with the topic being chosen in order to recognize the contributions that they have made to the study of the galactic evolution and in particular the early phases of the evolution of our Galaxy. The colloquium is sponsored by Observatoire de Paris, CFH telescope, and INSU-CNRS.

Scope of the colloquium The purpose is to bring together astronomers working in the field of the galaxies evolution (Milky Way and external galaxies), to review recent developments and constrain the models of the evolution of the galaxies using the properties of the old galactic components:

– the spherical halo,

– the thick disk,

– the bulge. We would like to encourage the participants to address in particular the work which is now in the process of being done with the arrival of the new generation of the 8m-10m telescopes. Several groups all over the world are actively involved in this type of research, and by combining different expertises it would be possible to achieve an optimal use of the very large telescopes. Exchange of information between different research groups as well as the promotion of a coordinated approach are essential for this purpose.

Outline of the scientific program

• Pregalactic evolution and nucleosynthesis

• Key features of stellar evolution

• Formation and evolution of galactic halos (Milky Way, External galaxies)

• The old disk component (Milky Way, External galaxies)

• Formation and evolution of galactic Bulges (Milky Way, External galaxies)

• Future Prospects

Scientific Organizing Committee Andersen J., Barbuy B., Beers T., Boesgaard A., Chmielevski Y., Friel E., Gratton R., Gustafsson B., Kunth D., Lelièvre G., Rich R.M., Spite F., Turon C.

Local Organizing Committee Baglin A., Briot D., Crifo F., Jablonka P., Spite M., Adam C., Bentolila C., Kervegant E., Pluet J.

Contact Monique Spite Observatoire de Paris-Meudon, F92195 Meudon cedex (France) FAX: +33-1-45 07 78 78

e-mail: collpm@obspm.fr

Web page See http://www.obspm.fr/collpm for further information about the colloquium, and registration.